OpenStudy (anonymous):

A circle has a radius of 5. What is the length of an arc that is subtended by a central angle of 55 degrees?

OpenStudy (anonymous):

9.59444444

OpenStudy (anonymous):

How come?

OpenStudy (abb0t):

Arc length = \(\sf \color{}{\frac{angle ~subtended~at~ centre}{360} \times 2 \pi r}\)

OpenStudy (anonymous):

is 4.8 correct?

OpenStudy (anonymous):

55/360 * 2*3.142*5

OpenStudy (the_fizicx99):

The answer is 4.7996554429844063365401496133437 .. rounding it would be what?

OpenStudy (ranga):

4.8 is correct.

OpenStudy (anonymous):

I am sorry...you are right...I divided by 180 and ended with wrong answer...